# Active Hypothesis Testing under Computational Budgets with Applications to GWAS and LLM

### v1

@@ 5.1 Myocardial Infarction GWAS
We apply our Active-Default framework to this task. Since small $p$-values are the signal of interest (an inverse signal), we use the utility function $u_{i}(x)=\log(1+1/(x+\epsilon))$ with $\epsilon=10^{-8}$. As the two GWAS were conducted on distinct cohorts, we assume the auxiliary and target $p$-values are independent and use the corresponding active $p$-value construction from (6). We include Random, Xu and Active-Xu for comparison.
Since the ground truth is unknown, we establish an oracle set of discoveries to serve as a benchmark, defined as the SNPs rejected by the BY procedure at $\alpha=0.1$ on the full set of $N$ MI $p$-values. We compare the performance of our Active-Default method against Random, Xu, and Active-Xu by measuring their ability to recover these oracle discoveries. For each method, we generate active $p$-values and apply the BY procedure to identify discoveries. Performance is quantified by efficiency, defined as the number of oracle discoveries recovered per MI $p$-value queried. We evaluate this efficiency as the budget, $n_{b}$, is varied as a fraction of the total number of SNPs, with the results summarized in Figure 7.
@@
The left panel of Figure 7 confirms that Active-Default, Active-Xu and the Random baseline precisely adhere to the specified budget, while Xu is unable to provide a budget guarantee. The right panel demonstrates the practical benefit of our approach. At every budget level, the Active-Default method is substantially more efficient than Random and Xu. This indicates that the HTN summary statistics, while not directly valid for inference, provide valuable information for prioritizing the analysis of MI associations, and our framework successfully exploits this information to maximize the return on the computational budget.
To provide external validation, we examine one of the top signals prioritized and discovered by our method, rs1333047. This SNP is located in the 9p21.3 locus, which is one of the most well-established and replicated risk loci for cardiovascular disease. A recent meta-analysis confirmed its strong association with coronary artery disease and MI (Paquette et al.,, 2017), suggesting that testing for this variant could be an important modifier of cardiovascular risk assessment.
@@ 5.2 Myocardial Infarction Complications
Our second real-data application also addresses myocardial infarction. We utilize a dataset from Golovenkin et al., (2020) that contains clinical information for 1,700 patients. The primary objective is to predict in-hospital mortality. The original dataset features a multi-class outcome with eight categories: survival and seven distinct causes of death. For our analysis, we simplify this into a testing problem: survival versus mortality, irrespective of the specific cause.
@@
To generate an auxiliary statistic, $P_{i}^{a}$, without this feature, we leverage an LLM (ChatGPT). We prompt the LLM to impute the ‘ZSN_A’ value for each patient based on their other clinical data. Using this completed dataset, we then compute a proxy conformal $p$-value, $P_{i}^{a}$, to serve as our auxiliary statistic.
@@
We then apply our Active-Default framework with a budget of $n_{b}=100$ and a significance level of $\alpha=0.1$ and compare it with Random, Xu, ALL, and Active-Xu baselines. The other settings are the same as in the previous subsection. Discoveries are identified using the BH procedure. While the theoretical validity of BH relies on the PRDS condition, we justify its use here on structural grounds. First, the underlying exact conformal $p$-values satisfy PRDS Bates et al., (2023). Second, although our active $p$-value is not a strictly monotonic transformation of the exact $p$-value, we expect it to remain highly positively correlated with it, thereby preserving the PRDS structure required for FDR control. The results are presented in Figure 8. While all five procedures successfully control the FDR below $\alpha=0.1$, they exhibit marked differences in budget adherence. The budgeted methods (Active-Default, Active-Xu, and Random) precisely respect the $n_{b}=100$ query limit, whereas the unconstrained Xu and ALL methods require significantly more computations. This highlights a clear trade-off between statistical power and computational cost. Although ALL and Xu achieve higher absolute power, our proposed Active-Default demonstrates the highest efficiency, delivering the greatest number of discoveries per query. The value of our global allocation scheme is further underscored by the comparison between Xu and Active-Xu; by enforcing the budget, Active-Xu achieves superior efficiency over its unconstrained counterpart. These findings collectively demonstrate that in resource-constrained settings where efficiency is paramount, our Active-Default framework provides a powerful and principled solution.
@@ 6 Discussion
In this work, we developed a general and theoretically grounded framework for active hypothesis testing under a global budget. Our method addresses the challenge of performing statistical inference when the computation of $p$-values or $e$-values is resource-intensive. By using a data-adaptive allocation scheme guided by auxiliary statistics, the framework produces a valid inferential outcome for every hypothesis while ensuring the expected number of expensive computations adheres to a pre-specified limit.
The practical implementation of our framework is guided by the choice of utility functions $\{u_{i}\}$ and the hyperparameter $\beta$. While our admissibility results (Propositions 1 and 2) show that no universally optimal choice exists, we have provided principle-driven guidance that yields effective performance in practice. These considerations also suggest several directions for future work. One direction is the development of data-driven methods for learning better utility functions. One could envision using a held-out calibration dataset to tune the form of $u_{i}(\cdot)$ to maximize a downstream objective, such as the number of discoveries, turning the selection process from a heuristic choice into a formal optimization problem.
Another significant extension would be to adapt our framework from a global budget in expectation to a strict, deterministic budget (i.e., $\sum C_{i}\leq n_{b}$ almost surely). This is a more challenging problem that would likely require techniques from sequential decision-making and online learning, where the decision to query a hypothesis is made adaptively based on the outcomes of previous queries. Finally, our framework could be extended to handle more complex structured inference problems, such as testing hypotheses on a graph or in a sequential, online setting where hypotheses arrive over time.

### v2

@@ 5.1 Myocardial Infarction GWAS
We apply our Active-Default framework to this task. Since small $p$-values are the signal of interest (an inverse signal), we use the utility function $1/(x+\epsilon)$ with $\epsilon=10^{-8}$ as the base utility function. As the two GWAS were conducted on distinct cohorts, we assume the auxiliary and target $p$-values are independent and use the corresponding active $p$-value construction from (6). We include Random, Xu and Active-Xu for comparison.
Since the ground truth is unknown, we establish an oracle set of discoveries to serve as a benchmark, defined as the SNPs rejected by the BY procedure at $\alpha=0.1$ on the full set of NN MI $p$-values. We compare the performance of our Active-Default method against Random, Xu, and Active-Xu by measuring their ability to recover these oracle discoveries. For each method, we generate active $p$-values and apply the BY procedure to identify discoveries. Performance is quantified by efficiency, defined as the number of oracle discoveries recovered per MI $p$-value queried. We evaluate this efficiency as the budget, $n_{b}$, is varied as a fraction of the total number of SNPs, with the results summarized in Figure 3.
@@
The left panel of Figure 3 confirms that Active-Default and Active-Xu precisely adhere to the specified budget, while Xu is unable to provide a budget guarantee. The right panel demonstrates the practical benefit of our approach. At every budget level, the Active-Default method is substantially more efficient than Random and Xu. This indicates that the HTN summary statistics, while not directly valid for inference, provide valuable information for prioritizing the analysis of MI associations, and our framework successfully exploits this information to maximize the return on the computational budget.
To provide external validation, we examine one of the top signals prioritized and discovered by our method, rs1333047. This SNP is located in the 9p21.3 locus, which is one of the most well-established and replicated risk loci for cardiovascular disease. A recent meta-analysis confirmed its strong association with coronary artery disease and MI (Paquette et al., 2017), suggesting that testing for this variant could be an important modifier of cardiovascular risk assessment.
@@ 5.2 Myocardial Infarction Complications
Our second real-data application also addresses myocardial infarction. We utilize a dataset from Golovenkin et al. (2020) that contains clinical information for 1,700 patients. The primary objective is to predict in-hospital mortality. The original dataset features a multi-class outcome with eight categories: survival and seven distinct causes of death. For our analysis, we simplify this into a testing problem: survival versus mortality, irrespective of the specific cause.
@@
To generate an auxiliary statistic, $P_{i}^{a}$, without this feature, we leverage a large language model, specifically Gemini 3.1 Pro. We prompt the LLM to impute the missing ‘ZSN_A’ value for each patient based on their other clinical data. Using this imputed dataset, we then compute a proxy conformal $p$-value, $P_{i}^{a}$, to serve as our auxiliary statistic. The complete prompt, the full conversation history, and the attached dataset are available at https://gemini.google.com/share/4922ddaad736.
@@
We then apply our Active-Default framework with a budget of $n_{b}=100$ and a significance level of $\alpha=0.1$ and compare it with Random, Xu, ALL, and Active-Xu baselines. As with the GWAS analysis, we treat the auxiliary statistic as inverse signals and use the base utility function $1/(x+\epsilon)$. The other settings are the same as in the previous subsection. Discoveries are identified using the BH procedure. While the theoretical validity of BH relies on the PRDS condition, we justify its use here on structural grounds. First, the underlying exact conformal $p$-values satisfy PRDS (Bates et al., 2023). Second, although our active $p$-value is not a strictly monotonic transformation of the exact $p$-value, we expect it to remain highly positively correlated with it, thereby preserving the PRDS structure required for FDR control. The results are presented in Figure 4. While all five procedures successfully control the FDR below $\alpha=0.1$, they exhibit marked differences in budget adherence. The budgeted methods (Active-Default, Active-Xu, and Random) precisely respect the $n_{b}=100$ query limit, whereas the unconstrained Xu and ALL methods require significantly more computations. This highlights a clear trade-off between statistical power and computational cost. Although ALL and Xu achieve higher absolute power, our proposed Active-Default demonstrates the highest efficiency, delivering the greatest number of discoveries per query. The value of our global allocation scheme is further underscored by the comparison between Xu and Active-Xu; by enforcing the budget, Active-Xu achieves superior efficiency over its unconstrained counterpart. These findings collectively demonstrate that in resource-constrained settings where efficiency is paramount, our Active-Default framework provides a powerful and principled solution.
@@ 6 Discussion
In this work, we developed a general and theoretically grounded framework for active hypothesis testing under a global budget. Our method addresses the challenge of performing statistical inference when the computation of $p$-values or $e$-values is resource-intensive. By using a data-adaptive allocation scheme guided by auxiliary statistics, the framework produces a valid inferential outcome for every hypothesis while ensuring that the exact number of expensive computations adheres to a pre-specified limit.
The practical implementation of our framework is guided by the choice of utility functions $\{u_{i}\}$ and the hyperparameter $\beta$. While our admissibility results (Propositions 1 and 2) show that no universally optimal choice exists, we have provided guidance that yields effective performance in practice. These considerations also suggest several directions for future work. One direction is the development of data-driven methods for learning better utility functions. One could envision using a held-out calibration dataset to tune the form of $u_{i}(\cdot)$ to maximize a downstream objective, such as the number of discoveries, turning the selection process from a heuristic choice into a formal optimization problem.
Another significant extension would be to handle more complex structured inference problems, such as testing hypotheses on a graph or in a sequential, online setting where hypotheses arrive over time. We discuss some preliminary ideas for the online setting in Supplement H.